# High Dynamic Range Novel View Synthesis with Single Exposure

### v1

@@ Abstract
High Dynamic Range Novel View Synthesis (HDR-NVS) aims to establish a 3D scene HDR model from Low Dynamic Range (LDR) imagery. Typically, multiple-exposure LDR images are employed to capture a wider range of brightness levels in a scene, as a single LDR image cannot represent both the brightest and darkest regions simultaneously. While effective, this multiple-exposure HDR-NVS approach has significant limitations, including susceptibility to motion artifacts (e.g., ghosting and blurring), high capture and storage costs. To overcome these challenges, we introduce, for the first time, the single-exposure HDR-NVS problem, where only single exposure LDR images are available during training. We further introduce a novel approach, Mono-HDR-3D, featuring two dedicated modules formulated by the LDR image formation principles, one for converting LDR colors to HDR counterparts, and the other for transforming HDR images to LDR format so that unsupervised learning is enabled in a closed loop. Designed as a meta-algorithm, our approach can be seamlessly integrated with existing NVS models. Extensive experiments show that Mono-HDR-3D significantly outperforms previous methods. Source code will be released.
@@ 1 Introduction
To address these issues, we propose a more deployable yet more challenging task, namely HDR-NVS with single-exposure LDR images, which eliminates the reliance on multiple exposures and thus avoids the aforementioned limitations. However, as illustrated in Fig. 1, single-exposure images frequently suffer from overexposure or underexposure, posing significant challenges for HDR-NVS. Furthermore, we present a novel HDR 3D scene modeling framework, Mono-HDR-3D, characterized by learning to approximate the underlying LDR image formation process of camera imaging. Specifically, we start with learning a LDR 3D scene model from single exposure LDR training images, followed by lifting the color space to HDR with a dedicated color transformation module in a per-channel manner. This from-LDR-to-HDR design is opposite to previous models since single exposure LDR images are insufficient for deriving an HDR model. We further introduce a closed-loop design by augmenting a process of converting HDR images to LDR images, allowing additional supervision even in case of no access to HDR ground-truth training data.
@@ 3.2 Mono-HDR-3D
Camera imaging mechanism
@@
Figure 3: Structure of our camera imaging aware LDR-to-HDR Color Converter (L2H-CC). $c_{i}^{l}$/$c_{i}^{h}$: LDR/HDR color; LO: Linear Operation, R: ReLU, SM: Sigmoid. $\odot$ and $\oplus$: Element-wise multiplication and addition.
L2H-CC
@@ 4 Experiments
Evaluation metrics. We employ the PSNR and SSIM as quantitative metrics. We utilize LPIPS as a perceptual metric, where lower values signify better perceptual quality. Note, for qualitative evaluation, we do not apply tone post-processing with third-party software which might introduce uncontrollable factors that distort the evaluation.
@@ 4.1 Quantitative evaluation
(II) Regarding the LDR NVS results, we observe a similar performance advantage with our models. This indicates that directly learning an HDR model from single-exposure LDR data, as competitors do, would be inferior due to the absence of multiple exposure observations. This also partly explains our stronger HDR NVS results, which somehow are dependent on the quality of the LDR input.
@@
Result analysis on real data While less important, Tab. 2 presents the quantitative results of LDR NVS on the real datasets, as there are no ground-truth HDR images available. The results are averaged across three distinct exposure times and encompass all scenes. It is evident that no method clearly stands out in synthesis quality.
@@ 4.3 Ablation Studies
Effect of closed-loop design
Following the above design based ablation, we further look into the effect of our closed-loop design with H2L-CC in the absence of HDR training data. The results in Tab. 4 indicate that this design brings a positive impact of 0.38 dB increase in PSNR for HDR NVS.

### v2

@@ Abstract
High Dynamic Range Novel View Synthesis (HDR-NVS) aims to establish a 3D scene HDR model from Low Dynamic Range (LDR) imagery. Typically, multiple-exposure LDR images are employed to capture a wider range of brightness levels in a scene, as a single LDR image cannot represent both the brightest and darkest regions simultaneously. While effective, this multiple-exposure HDR-NVS approach has significant limitations, including susceptibility to motion artifacts (e.g., ghosting and blurring), high capture and storage costs. To overcome these challenges, we introduce, for the first time, the single-exposure HDR-NVS problem, where only single exposure LDR images are available during training. We further introduce a novel approach, Mono-HDR-3D, featuring two dedicated modules formulated by the LDR image formation principles, one for converting LDR colors to HDR counterparts and the other for transforming HDR images to LDR format so that unsupervised learning is enabled in a closed loop. Designed as a meta-algorithm, our approach can be seamlessly integrated with existing NVS models. Extensive experiments show that Mono-HDR-3D significantly outperforms previous methods. Source code is released at https://github.com/prinasi/Mono-HDR-3D.
@@ 1 Introduction
To address these issues, we propose a more deployable yet more challenging task, namely HDR-NVS with single-exposure LDR images, which eliminates the reliance on multiple exposures and thus avoids the aforementioned limitations. However, as illustrated in Fig. 1, single-exposure images frequently suffer from overexposure or underexposure, posing significant challenges for HDR-NVS. Furthermore, we present a novel HDR 3D scene modeling framework, Mono-HDR-3D, characterized by learning to approximate the underlying LDR image formation process of camera imaging. Specifically, we start with learning an LDR 3D scene model from single-exposure LDR training images, followed by lifting the color space to HDR with a dedicated color transformation module in a per-channel manner. This from-LDR-to-HDR design is opposite to previous models since single-exposure LDR images are insufficient for deriving an HDR model. We further introduce a closed-loop design by augmenting a process of converting HDR images to LDR images, allowing additional supervision even in the case of no access to HDR ground-truth training data.
@@ 3.2 Mono-HDR-3D
Camera imaging mechanism.
@@
Figure 3: Structure of our camera imaging aware LDR-to-HDR Color Converter (L2H-CC). $c_{i}^{l}$/$c_{i}^{h}$: LDR/HDR color; LO: Linear Operation, R: ReLU, SP: Softplus. $\odot$ and $\oplus$: Element-wise multiplication and addition.
L2H-CC.
@@ 4 Experiments
Evaluation metrics. We employ the PSNR and SSIM as quantitative metrics. We utilize LPIPS as a perceptual metric, where lower values signify better perceptual quality. Similar to HDR-GS (Cai et al., 2024), we also quantitatively evaluate the rendered HDR images in the tone-mapped domain and qualitatively show HDR results tone-mapped by Photomatrix Pro (HDRsoft Team, 2025).
@@ 4.1 Quantitative evaluation
(II) Regarding the LDR NVS results, we observe a similar performance advantage with our models. This indicates that directly learning an HDR model from single-exposure LDR data, as competitors do, would be inferior due to the absence of multiple exposure observations. This also partly explains our stronger HDR NVS results, which somehow are dependent on the quality of the LDR output.
@@
Result analysis on real data. While less important, Tab. 2 presents the quantitative results of LDR NVS on the real datasets, as there are no ground-truth HDR images available. The results are averaged across three distinct exposure times and encompass all scenes. It is evident that no method clearly stands out in synthesis quality.
@@ 4.3 Ablation studies
Effect of closed-loop design.
Following the above design based ablation, we further look into the effect of our closed-loop design with H2L-CC with HDR training data. The results in Tab. 4 indicate that this design brings a positive impact of 0.38 dB increase in PSNR for HDR NVS, demonstrating that the closed-loop framework contributes significantly to enhancing the quality of the reconstructed HDR images.
Loss contributions.
Based on our proposed Mono-HDR-GS, we conduct systematic ablation studies to evaluate the contribution of each loss component in Eq. (9). The results in Tab. 5 reveal three key observations: (I) Row 2 vs. Row 7: HDR loss $L_{\text{hdr}}$ serves as the foundational component for HDR-NVS performance. When used alone, Mono-HDR-GS achieves moderate metrics (33.93 dB PSNR, 0.925 SSIM), while its absence leads to severe degradation (e.g., Row 5 shows 13.50 dB PSNR without $L_{\text{hdr}}$). This validates its critical role in reconstructing high dynamic range scenes. (II) Row 4 vs. Row 6: LDR loss $L_{\text{ldr}}$ provides essential regularization for scene modeling. When combined with $L_{\text{hdr}}$, it improves PSNR by +4.26 dB (Row 2 vs. 4) and maintains structural fidelity. Notably, $L_{\text{ldr}}$ alone fails to train (Row 1), but acts as a complementary constraint when paired with HDR-aware objectives. (III) Row 4 vs. 7: Closed-loop loss $L_{\text{h2l}}$ enhances photometric consistency across exposure domains. Its integration elevates PSNR by +0.38 dB while reducing perceptual distance. This demonstrates that bidirectional optimization between HDR and LDR spaces refines both radiance estimation and tone mapping.
Table 5: Ablation analysis of different loss components on the synthetic datasets. Results are averaged across exposures and scenes.
Row Loss HDR result
[…]
1 $L_{\text{ldr}}$
2 $L_{\text{hdr}}$ 33.93 0.925 0.050
3 $L_{\text{h2l}}$ 11.87 0.504 0.371
4 $L_{\text{ldr}}+L_{\text{hdr}}$ 38.19 0.974 0.015
5 $L_{\text{ldr}}+L_{\text{h2l}}$ 13.50 0.507 0.359
6 $L_{\text{hdr}}+L_{\text{h2l}}$ 33.58 0.934 0.058
7 $L_{\text{ldr}}+L_{\text{hdr}}+L_{\text{h2l}}$ 38.57 0.975 0.012
Performance of different LDR / HDR ratios.
To evaluate the robustness of our Mono-HDR-GS framework under varying data availability, we conduct experiments with LDR / HDR ratios ranging from 0/1 (pure HDR) to 5/1 (dominant LDR), as shown in Tab. 6. The results demonstrate:
(I) Data Efficiency: Our model maintains strong performance even with sparse HDR data. For instance, at 5/1 ratio, Mono-HDR-GS retains 92.6% of its peak PSNR (35.51 dB vs. 38.57 dB at 1/1), suggesting effective knowledge transfer from LDR supervision.
(II) LDR Supervision: When HDR data becomes extremely scarce (5/1 vs. 0/1), the PSNR degradation of Mono-HDR-GS (from 35.51 dB to 33.93 dB) is significantly smaller than HDR-GS (from 34.89 dB to 33.46 dB). This confirms that LDR supervision provides a more robust geometric prior for HDR reconstruction.
(III) HDR Criticality: Pure HDR supervision (0/1) outperforms pure LDR supervision (1/0) by +23.4 dB PSNR, validating the irreplaceable role of HDR data in capturing radiance information. This finding emphasizes HDR’s fundamental importance for high-quality HDR novel view synthesis.
(IV) Overall Superiority: Across all ratios, Mono-HDR-GS consistently outperforms HDR-GS with statistically significant margins. At equal computational cost, our Mono-HDR-GS achieves up to +3.27 dB PSNR improvement (1/1 ratio) and maintains superior structural fidelity (SSIM 0.975 vs. 0.965). Notably, even with 100% LDR data (1/0 ratio), our method generates marginally better results than HDR-GS trained solely on HDR images, demonstrating the effectiveness of closed-loop design.
Table 6: Ablation studies of different ratio of LDR / HDR images.
Method LDR / HDR HDR result
[…]
HDR-GS 1/1 35.30 0.965 0.030
Mono-HDR-GS 38.57 0.975 0.012
HDR-GS 2/1 35.26 0.963 0.033
Mono-HDR-GS 37.97 0.975 0.013
HDR-GS 3/1 35.16 0.958 0.035
Mono-HDR-GS 37.53 0.974 0.014
HDR-GS 5/1 34.89 0.961 0.027
Mono-HDR-GS 35.51 0.963 0.023
HDR-GS 0/1 33.46 0.936 0.075
Mono-HDR-GS 33.93 0.925 0.050
HDR-GS 1/0 10.51 0.503 0.350
Mono-HDR-GS 13.50 0.507 0.359
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the Key Research and Development Plan of Jiangsu Province (Industry Foresight and Key Core Technology Project) under Grant No. BE2023008-2.